OpenStudy (anonymous):

Which statement is true for the circle? Its center is at (3/2,6) and the radius is 16/9 Its center is at (3/2,-6) and the radius is 4/3 Its center is at (-3/2,-6) and the radius is 4/3 Its center is at (-3/2,6) and the radius is 16/9

myininaya (myininaya):

The equation of the circle is missing

OpenStudy (anonymous):

what does the circle look like we need the equation

myininaya (myininaya):

\[(x-h)^2+(y-k)^2=r^2 \] (h,k) is center and r is radius

myininaya (myininaya):

\[(x+\frac{3}{2})^2+(y+6)^2=\frac{16}{9}\] So what is -h= -k= r^2=

myininaya (myininaya):

Torie, compare those equations what is in the place of -h in your equation?

myininaya (myininaya):